OpenStudy (anonymous):

5.5% tax on $20 is $__

OpenStudy (anonymous):

5.5/100 * 20 = x 55/1000 * 20 = x 55*2/100 = x 110/100 = x 11/10 = x 1.1 = x so, 5.5% tax on $20 is $1.1